# Stopping Rules for Stochastic Gradient Descent via Anytime-Valid Confidence Sequences

### v2

@@ 2 Anytime-Valid Confidence Sequences
$B_{k}(v):=\frac{\log(1/(\alpha w_{k}))}{\lambda_{k}}+2\lambda_{k}v=\frac{e^{k/2}}{\lambda_{0}}\big(L_{\alpha}+2\log(k+1)\big)+2\lambda_{0}e^{-k/2}v.$
@@
    $e^{k(v)/2}$ is within a constant factor of
@@
which immediately implies that $e^{k(v)/2}\leq z_{v}\leq e^{(k(v)+1)/2}$. Assumption 2.1 and Lemma A.3 (note that $L_{\alpha}\geq 1$ holds since $\alpha\leq 0.6$) guarantee that $k(v)\geq 0$. Therefore, for this choice, we have that
$\frac{1}{\sqrt{e}}\,z_{v}\leq e^{k(v)/2}\leq z_{v},\qquad\log(k(v)+1)\leq c_{1}L_{v},$ (21)
@@
$e^{k(v)/2}\leq z_{v}=\sqrt{\frac{v}{L_{\alpha}+L_{v}}},\qquad e^{-k(v)/2}\leq\sqrt{e}\sqrt{\frac{L_{\alpha}+L_{v}}{v}}.$
@@
$\frac{e^{k(v)/2}}{\lambda_{0}}D_{v}\leq\frac{1}{\lambda_{0}}\sqrt{\frac{v}{L_{\alpha}+L_{v}}}\,c_{2}\big(L_{\alpha}+L_{v}\big)=\frac{c_{2}}{\lambda_{0}}\sqrt{v\big(L_{\alpha}+L_{v}\big)},$
@@
$2\lambda_{0}e^{-k(v)/2}v\leq 2\lambda_{0}\sqrt{e}\sqrt{\frac{L_{\alpha}+L_{v}}{v}}\,v=2\sqrt{e}\lambda_{0}\sqrt{v\big(L_{\alpha}+L_{v}\big)}.$
Summing these two bounds and recalling that $\lambda_{0}=\sqrt{{c_{2}}/({2\sqrt{e}})}$, we obtain
$B_{k(v)}(v)\leq 2\sqrt{2\sqrt{e}c_{2}}\sqrt{v\big(L_{\alpha}+L_{v}\big)}\leq c_{3}\sqrt{v\big(L_{\alpha}+L_{v}\big)},$
@@
$\inf_{k\geq 0}B_{k}(v)\leq C\sqrt{v\big(L_{\alpha}+\log\log(\mathrm{e}+v)\big)}=C\sqrt{v\big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+v)\big)},$
where $c\coloneqq\log\tfrac{\pi^{2}}{6}$.
@@
$\bar{X}_{t}\leq C\sqrt{\Sigma_{t}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}\qquad\forall t\geq 1,$
@@
Let $X_{t}$ be defined as in (12), and let $\sigma_{t}^{2}$ and $\Sigma_{t}^{2}$ be as in (8). Then, for any $\alpha\in(0,0.6)$, with probability at least $1-\alpha$, the following holds simultaneously for all $t\geq 1$:
$\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)\;\leq\;\frac{1}{2}\Bigg(C\,\sqrt{\Sigma_{t}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg),$ (22)
@@
From Lemma 2.7 we know that for any $\alpha\in(0,0.6)$, there exists an event $\mathcal{E}_{\alpha}$ with $\mathbb{P}(\mathcal{E}_{\alpha})\geq 1-\alpha$ such that, on $\mathcal{E}_{\alpha}$,
$\forall t\geq 1:\qquad\sum_{s=1}^{t}X_{s}\leq C\sqrt{\Sigma_{t}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}.$ (23)
@@
$2\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)\leq C\sqrt{\Sigma_{t}^{2}\Big(\log\tfrac{1}{\alpha}+c+\log\log(\mathrm{e}+\Sigma_{t}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}.$
@@ 2.2 Clarifying remarks and intuition
$\bar{X}_{t}<\frac{\log(1/\alpha)}{\lambda}+2\lambda\Sigma_{t}^{2}\qquad\forall t\geq 1.$ (26)
@@
$B(\lambda,t):=\frac{\log(1/\alpha)}{\lambda}+2\lambda\Sigma_{t}^{2},\qquad\lambda>0.$ (27)
@@
$\lambda^{\star}(t)=\sqrt{\frac{\log(1/\alpha)}{2\Sigma_{t}^{2}}},$ (28)
@@
$B\big(\lambda^{\star}(t),t\big)=2\sqrt{2}\sqrt{\Sigma_{t}^{2}\log\frac{1}{\alpha}}.$ (29)
Thus, for a fixed time $t$, a single exponential supermartingale produces an upper bound with the scaling $\sqrt{\Sigma_{t}^{2}\log(1/\alpha)}$. This matches the leading term that appears in Theorem 2.2, but without the additional iterated-logarithm factor that ultimately arises from restricting $\lambda$ to lie on a fixed grid.
The problem arises from the fact that $\lambda^{\star}(t)$ in (28) depends on $\Sigma_{t}^{2}$ and is therefore $\mathcal{F}_{t}$-measurable: it is chosen after observing the data up to time $t$. Substituting this data-dependent $\lambda^{\star}(t)$ back into $E_{t}(\lambda)$ destroys the supermartingale structure, so Ville’s inequality cannot be applied time-uniformly. With a single exponential process one faces a choice: either fix $\lambda$ in advance (yielding time-uniform but suboptimal bounds for many $t$), or optimize $\lambda$ at each $t$ (yielding sharp fixed-time bounds but no optional-stopping guarantee).
The mixture construction in Lemma 2.7 reconciles these goals by fixing a countable family of parameters a priori and combining the corresponding exponential supermartingales. This “method of mixtures” (also called pseudo-maximization) goes back to Robbins and Siegmund’s classical work on sequential boundary crossing and is developed systematically in the theory of self-normalized processes (see [7]). Modern confidence-sequence constructions build on this principle in various forms, including explicit mixture and stitching frameworks such as the sub-$\psi$ approach in [15], as well as betting- and e-process-based methods that aggregate families of nonnegative supermartingales; see, [29]. $\clubsuit$
@@
The mixture construction in Lemma 2.7 hinges on choosing a geometric grid of tuning parameters $\lambda_{k}=\lambda_{0}e^{-k/2}$. Here we provide intuition for this choice by examining how the envelope $B_{k}(\Sigma_{t}^{2})$ is minimized, and how the geometry of the grid matches the natural scale of the problem. Fix $v>0$ and recall the upper envelope
$B_{k}(v):=\frac{\log(1/(\alpha w_{k}))}{\lambda_{k}}+2\lambda_{k}v,\qquad\lambda_{k}=\lambda_{0}e^{-k/2},\quad w_{k}=\frac{6}{\pi^{2}(k+1)^{2}}.$
Momentarily regard $k$ as a continuous variable and write $z=e^{k/2}>0$. Then
@@
Heuristically, since $D(z)\approx L_{\alpha}+\log\log(\mathrm{e}+z)\approx L_{\alpha}+\log\log(\mathrm{e}+v)$ at the relevant scale, this suggests the leading-order behavior
$\inf_{k\geq 0}B_{k}(v)\;\lesssim\;\sqrt{v\Big(L_{\alpha}+\log\log(\mathrm{e}+v)\Big)}.$
@@
The slowly varying term $\log(k+1)$ introduces an iterated-logarithm correction, where the $\log\log(\mathrm{e}+v)$ factor arises precisely from the slow growth of $\log(k+1)$ on this geometric grid. In particular, for a fixed time $t$, the optimal value of $\lambda$ is of order $v^{-1/2}$, where $v=\Sigma_{t}^{2}$ controls the variance scale of the martingale sum. Since $v$ can vary over many orders of magnitude as $t$ increases, a geometric grid of the form $\lambda_{k}=\lambda_{0}e^{-k/2}$ is a natural discretization: consecutive grid points differ by a constant factor, ensuring that at least one $\lambda_{k}$ lies within a constant factor of the ideal $\lambda^{\star}\approx v^{-1/2}$, for every $t$.
@@
In short, the grid is chosen so that: (i) it spans the relevant scales of $\Sigma_{t}^{2}$ via exponentially spaced search points, and (ii) at least one $\lambda_{k}$ is guaranteed to be within a constant factor of the ideal minimizer, thereby achieving near-optimal behavior across all times $t$. This “scale-matching” intuition for geometric discretizations is common in mixture and stitching arguments for martingale concentration; see, for example, the heuristic discussion surrounding Theorem 1 in [15] and the broader treatment of pseudo-maximization via mixtures in self-normalized processes [7]. $\clubsuit$
Remark 2.11 (Why no $\log\log t$ term appears).
@@
A recurring feature in time-uniform martingale concentration is the appearance of iterated logarithm corrections, such as $\log\log t$; see, for example, classical boundary-crossing analyses of Robbins and Siegmund [26] and modern confidence-sequence treatments such as [15]. At first glance, one might therefore expect a confidence sequence for stochastic optimization to include a term of the form $\sqrt{\Sigma_{t}^{2}\log\log t}$ rather than the $\sqrt{\Sigma_{t}^{2}\log\log(\mathrm{e}+\Sigma_{t}^{2})}$ term appearing in Theorem 2.2. The difference comes from the square-summable stepsize assumption in Assumption 1.2 and the resulting definition of $\Sigma_{t}^{2}=\sum_{s=1}^{t}\sigma^{2}\eta_{s}^{2}Z_{s}$. In most sequential testing settings, $\Sigma_{t}^{2}$ grows linearly with $t$, so controlling $k$ in Step (iv) of Lemma 2.7 ultimately leads to iterated-logarithm terms in $\log t$. Here, however, $\{\eta_{s}\}_{s\geq 1}$ is square summable, so $\Sigma_{t}^{2}$ is bounded and converges to
$\Sigma_{\infty}^{2}:=\sum_{s=1}^{\infty}\sigma^{2}\eta_{s}^{2}Z_{s}\;<\;\infty\qquad\text{almost surely},$
under Assumption 1.2. Consequently, the relevant “scale” for the mixture optimization is the value of $\Sigma_{t}^{2}$ itself, not the time index $t$. This is why the iterated-logarithm term takes the form $\log\log(\mathrm{e}+\Sigma_{t}^{2})$ rather than $\log\log t$: the penalty adapts to the variance process instead of time. Because $\sum_{t\geq 1}\eta_{t}^{2}<\infty$ ensures that $\Sigma_{t}^{2}$ is almost surely bounded, this “variance-based iterated log” is substantially smaller than the classical “time-based iterated log” that arises when $\Sigma_{t}^{2}$ grows linearly with $t$.
@@
This phenomenon is well known in the martingale boundary literature: confidence sequences expressed in terms of data-dependent variance proxies (such as predictable quadratic variation, conditional variance processes, or self-normalized sums) adapt to the realized variability of the process and can be substantially tighter when the variance grows slowly or remains bounded, compared to bounds written purely in terms of time. See, for example, [7] and the discussion in [15, Section 3] on expressing time-uniform bounds in terms of intrinsic variance processes rather than the time index. $\clubsuit$

### v3

@@ 2 Anytime-Valid Confidence Sequences
$B_{k}(v):=\frac{\log(1/(\alpha w_{k}))}{\lambda_{k}}+2\lambda_{k}v\leq\frac{\mathrm{e}^{k/2}}{\lambda_{0}}\big(L_{\alpha}+2\log(k+1)\big)+2\lambda_{0}\mathrm{e}^{-k/2}v.$
@@
    $\mathrm{e}^{k(v)/2}$ is within a constant factor of
@@
which immediately implies that $\mathrm{e}^{k(v)/2}\leq z_{v}\leq\mathrm{e}^{(k(v)+1)/2}$. Since $\Sigma_{t,\mathrm{eff}}^{2}\geq 2(\log\tfrac{2}{\alpha}+1)$ by definition, Lemma A.3 guarantees that $k(v)\geq 0$. For the lemma, we note that $L_{\alpha}\geq 1$ holds since $\alpha\leq\tfrac{2}{\mathrm{e}}$. Therefore, for this choice, we have that
$\frac{1}{\sqrt{\mathrm{e}}}\,z_{v}\leq\mathrm{e}^{k(v)/2}\leq z_{v},\qquad\log(k(v)+1)\leq c_{1}L_{v},$ (22)
@@
$\mathrm{e}^{k(v)/2}\leq z_{v}=\sqrt{\frac{v}{L_{\alpha}+L_{v}}},\qquad\mathrm{e}^{-k(v)/2}\leq\sqrt{\mathrm{e}}\sqrt{\frac{L_{\alpha}+L_{v}}{v}}.$
@@
$\frac{\mathrm{e}^{k(v)/2}}{\lambda_{0}}D_{v}\leq\frac{1}{\lambda_{0}}\sqrt{\frac{v}{L_{\alpha}+L_{v}}}\,c_{2}\big(L_{\alpha}+L_{v}\big)=\frac{c_{2}}{\lambda_{0}}\sqrt{v\big(L_{\alpha}+L_{v}\big)},$
@@
$2\lambda_{0}\mathrm{e}^{-k(v)/2}v\leq 2\lambda_{0}\sqrt{\mathrm{e}}\sqrt{\frac{L_{\alpha}+L_{v}}{v}}\,v=2\sqrt{\mathrm{e}}\lambda_{0}\sqrt{v\big(L_{\alpha}+L_{v}\big)}.$
Summing these two bounds and recalling that $\lambda_{0}=\sqrt{{c_{2}}/({2\sqrt{\mathrm{e}}})}$, we obtain
$B_{k(v)}(v)\leq 2\sqrt{2\sqrt{\mathrm{e}}c_{2}}\sqrt{v\big(L_{\alpha}+L_{v}\big)}\leq c_{3}\sqrt{v\big(L_{\alpha}+L_{v}\big)},$
@@
$\inf_{k\geq 0}B_{k}(v)\leq C\sqrt{v\big(L_{\alpha}+\log\log(\mathrm{e}+v)\big)}=C\sqrt{v\big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+v)\big)}.$
@@
$\bar{X}_{t}\leq C\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}\qquad\forall t\geq 1,$
@@
For any $\alpha\in(0,\tfrac{2}{\mathrm{e}})$, with probability at least $1-\alpha$, the following holds simultaneously for all $t\geq 1$:
$\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)\;\leq\;\frac{1}{2}\Bigg(C\,\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}\Bigg),$ (23)
@@
From Lemma 2.6 we know that for any $\alpha\in(0,\tfrac{2}{\mathrm{e}})$, there exists an event $\mathcal{E}_{\alpha}$ with $\mathbb{P}(\mathcal{E}_{\alpha})\geq 1-\alpha$ such that, on $\mathcal{E}_{\alpha}$,
$\forall t\geq 1:\qquad\sum_{s=1}^{t}X_{s}\leq C\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}.$ (24)
@@
$2\sum_{s=1}^{t}\eta_{s}\big(f(x_{s})-f(x^{\star})\big)\leq C\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})\Big)}+Z_{1}+\sum_{s=1}^{t}\eta_{s}^{2}\|g_{s}\|^{2}.$
@@ 2.2 Clarifying remarks and intuition
$\bar{X}_{t}<\frac{\log(1/\alpha)}{\lambda}+2\lambda\Sigma_{t}^{2}\leq\frac{\log(1/\alpha)}{\lambda}+2\lambda\Sigma_{t,\mathrm{eff}}^{2}\qquad\forall t\geq 1.$ (27)
@@
$B(\lambda,t):=\frac{\log(1/\alpha)}{\lambda}+2\lambda\Sigma_{t,\mathrm{eff}}^{2},\qquad\lambda>0.$ (28)
@@
$\lambda^{\star}(t)=\sqrt{\frac{\log(1/\alpha)}{2\Sigma_{t,\mathrm{eff}}^{2}}},$ (29)
@@
$B\big(\lambda^{\star}(t),t\big)=2\sqrt{2}\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\log\frac{1}{\alpha}}\leq 2\sqrt{2}\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\log\frac{2}{\alpha}}.$ (30)
Thus, for a fixed time $t$, a single exponential supermartingale produces an upper bound with the scaling $\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\log(2/\alpha)}$. This matches the leading term that appears in Theorem 2.1, but without the additional iterated-logarithm factor that arises when enforcing time-uniform validity.
The problem arises from the fact that $\lambda^{\star}(t)$ in (29) depends on $\Sigma_{t,\mathrm{eff}}^{2}$ and is therefore $\mathcal{F}_{t}$-measurable: it is chosen after observing the data up to time $t$. Substituting this data-dependent $\lambda^{\star}(t)$ back into $E_{t}(\lambda)$ destroys the supermartingale structure, so Ville’s inequality cannot be applied time-uniformly. With a single exponential process one faces a choice: either fix $\lambda$ in advance (yielding time-uniform but suboptimal bounds for many $t$), or optimize $\lambda$ at each $t$ (yielding sharp fixed-time bounds but no optional-stopping guarantee).
The mixture construction in Lemma 2.6 reconciles these goals by fixing a countable family of parameters a priori and combining the corresponding exponential supermartingales. This “method of mixtures” (also called pseudo-maximization) goes back to Robbins and Siegmund’s classical work on sequential boundary crossing and is developed systematically in the theory of self-normalized processes (see [de2009self]). Modern confidence-sequence constructions build on this principle in various forms, including explicit mixture and stitching frameworks such as the sub-$\psi$ approach in [howard2021time], as well as betting- and e-process-based methods that aggregate families of nonnegative supermartingales; see [waudby2024estimating]. $\clubsuit$
@@
The mixture construction in Lemma 2.6 hinges on choosing a geometric grid of tuning parameters $\lambda_{k}=\lambda_{0}\mathrm{e}^{-k/2}$. Here we provide intuition for this choice by examining how the envelope $B_{k}(\Sigma_{t,\mathrm{eff}}^{2})$ is minimized, and how the geometry of the grid matches the natural scale of the problem. Let $v\coloneqq\Sigma_{t,\mathrm{eff}}^{2}$ and recall the upper envelope
$B_{k}(v):=\frac{\log(1/(\alpha w_{k}))}{\lambda_{k}}+2\lambda_{k}v,\qquad\lambda_{k}=\lambda_{0}\mathrm{e}^{-k/2},\quad w_{k}=\frac{6}{\pi^{2}(k+1)^{2}}.$
Momentarily regard $k$ as a continuous variable and write $z=\mathrm{e}^{k/2}>0$. Then
@@
Heuristically, since $D(z)\approx\log\tfrac{\pi^{2}}{6\alpha}+\log\log(\mathrm{e}+z)$, which at the relevant scale is equivalent up to constants to $\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+v)$, this suggests the leading-order behavior
$\inf_{k\geq 0}B_{k}(v)\;\lesssim\;\sqrt{v\Big(\log\tfrac{2}{\alpha}+\log\log(\mathrm{e}+v)\Big)}.$
@@
The slowly varying term $\log(k+1)$ introduces an iterated-logarithm correction, where the $\log\log(\mathrm{e}+v)$ factor arises precisely from the slow growth of $\log(k+1)$ on this geometric grid. In particular, for a fixed time $t$, the optimal value of $\lambda$ is of order $v^{-1/2}$, where $v=\Sigma_{t,\mathrm{eff}}^{2}$ controls the variance scale of the martingale sum. Since the scale $\Sigma_{t,\mathrm{eff}}^{2}$ is not known in advance and may vary with the trajectory up to the a priori bound $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$, a geometric grid of the form $\lambda_{k}=\lambda_{0}\mathrm{e}^{-k/2}$, with $\lambda_{0}$ chosen as a function of $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}$, is a natural discretization: consecutive grid points differ by a constant factor, ensuring that for every $t$ there exists a $\lambda_{k}$ within a constant factor of the ideal choice $\lambda^{\star}\approx\Sigma_{t,\mathrm{eff}}^{-1}$.
@@
In short, the grid is chosen so that: (i) it spans the relevant scales of $\Sigma_{t,\mathrm{eff}}^{2}$ via exponentially spaced search points, and (ii) at least one $\lambda_{k}$ is guaranteed to be within a constant factor of the ideal minimizer, thereby achieving near-optimal behavior across all times $t$. This “scale-matching” intuition for geometric discretizations is common in mixture and stitching arguments for martingale concentration; see, for example, the heuristic discussion surrounding Theorem 1 in [howard2021time] and the broader treatment of pseudo-maximization via mixtures in self-normalized processes [de2009self]. $\clubsuit$
Remark 2.10 (Why no $\log\log(\mathrm{e}+t)$ term appears).
@@
A recurring feature in time-uniform martingale concentration is the appearance of iterated logarithm corrections, such as $\log\log(\mathrm{e}+t)$; see, for example, classical boundary-crossing analyses of Robbins and Siegmund [robbins1970boundary] and modern confidence-sequence treatments such as [howard2021time]. At first glance, one might therefore expect a confidence sequence for stochastic optimization to include a term of the form $\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\log\log(\mathrm{e}+t)}$ rather than the $\sqrt{\Sigma_{t,\mathrm{eff}}^{2}\log\log(\mathrm{e}+\Sigma_{t,\mathrm{eff}}^{2})}$ term appearing in Theorem 2.1. The difference comes from the square-summable stepsize condition in Assumption 1.2 and the resulting definition of $\Sigma_{t,\mathrm{eff}}^{2}$. In most sequential testing settings, $\Sigma_{t,\mathrm{eff}}^{2}$ grows linearly with $t$, so controlling $k$ in Step (iv) of Lemma 2.6 ultimately leads to iterated-logarithm terms in $\log(\mathrm{e}+t)$. Here, however, $\{\eta_{t}\}_{t\geq 1}$ is square summable, so $\Sigma_{t,\mathrm{eff}}^{2}$ is uniformly bounded by $\widetilde{\Sigma}_{\infty,\mathrm{eff}}^{2}<\infty$. Consequently, the relevant “scale” for the confidence sequence is the variance process itself rather than the time index. This variance-based iterated logarithm therefore remains bounded and is substantially smaller than the classical time-based correction that arises when the variance grows linearly with $t$.
@@
This phenomenon is well known in the martingale boundary literature: confidence sequences expressed in terms of data-dependent variance proxies (such as predictable quadratic variation, conditional variance processes, or self-normalized sums) adapt to the realized variability of the process and can be substantially tighter when the variance grows slowly or remains bounded, compared to bounds written purely in terms of time. See, for example, [de2009self] and the discussion in [howard2021time, Section 3] on expressing time-uniform bounds in terms of intrinsic variance processes rather than the time index. $\clubsuit$